

********************************************************************
UARS WEEKLY STATUS REPORT - 14 JAN 2000
********************************************************************
***********
SOLSTICE
***********
We continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations.
SOLSTICE operates normally, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible.
***********
SUSIM
***********
SUSIM continues to gather a minimum set of daily solar scans scheduled
for periods of predicted telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM spectrometers continue to be active under PMON control*. While
under PMON control, the PEM spectrometers computed energy deposition from
electrons and protons measured in-situ by the particle spectrometers.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
The MPG participated in the January monthly UARS Mission
Status Review meeting on 12 January. Topics discussed
during this meeting included a discussion of the options
available for science data collection with the degraded
tape recorder. The recommendation from the meeting was
to continue to depend on real-time data via two TDRSS, and
to supplement this data coverage with tape recorded data.
The tape recorder will be programmed via stored commands
to capture all HALOE sunrise and sunset events that would
otherwise be missed during real-time contacts.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
13 days GPI (4.05) PROD (CLAES L3B) (456R-452C-4F=0L)
27 days WIN (5.11A) PROD (L1-L3) (520R-280C-11F=229L)
1 day SEN (5.0) PROD (1992) (1R-1C=0L)
9 days SUS (7.2) PROD (L1-L3) (TEST) (10R-9C=1L)
The CDHF completed processing GPI V4.05 software against all of the
CLAES V7.0 3AL data.
The CDHF is processing SENSOR V5.0 software against 1992 data as requested
by FDF.
The CDHF is testing SUSIM V7.2 Formal Software Delivery (Patch).
CDHF REHOST:
- Testing HRDI V4.4 on the Rehost System.
- Testing SOLSTICE V4.62 on the Rehost System.
- Plans to test WINDII V5.11A on the Rehost System.
- Rehost procedures/work instructions continued to be developed and/or
reviewed.
- The performance of Alpha CDHFA, CDHFB, and UCSS shutdown/ initialization
has been turned over from PAC to Operations personnel.
***********
FOT
***********
REPORTING PERIOD
This report covers 01 January, 2000 (Orbit 45394, GMT Day 001) through 07
January 2000 (Orbit 45499, GMT Day 007).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 003/2317 45439 Off
The beta angle ranged between the angles of 26.8 degrees to 33.2 degrees for
this report period. The beta angle is now decreasing toward a minimum angle of
+1.0 degrees on 06 January 1999 (DOY 006).
The clock error ranged from -6.2 to 1.9 msec during this report period. Clock
rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 01
% ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 02
% HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 01
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/358 - 2287.496855
Transponder B: 99/365 - 2287.499111
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/006 show the observatory data loss to be 560 hours, 29
minutes, 53.881 seconds (an increase of 81 hours 20 minutes 21.024 seconds
since the last report period). This is a 0.7713 percent data loss which equals
a 99.2287 percent data capture for the mission.
The increase of 81 hours 20 minutes 21.024 seconds occurred over seven (7)
days:
10 hours 26 minutes 12.992 seconds on Day 365 due to a S/C TR problem.
10 hours 47 minutes 35.616 seconds on Day 001 due to a S/C TR problem.
10 hours 37 minutes 26.784 seconds on Day 002 due to a S/C TR problem.
10 hours 59 minutes 12.608 seconds on Day 003 due to a S/C TR problem.
13 hours 09 minutes 55.968 seconds on Day 004 due to a S/C TR problem.
12 hours 43 minutes 32.704 seconds on Day 005 due to a S/C TR problem.
12 hours 36 minutes 24.352 seconds on Day 006 due to a S/C TR problem.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The Yaw maneuver went as planned and we are now flying in reverse. This
maneuver occured on January 6, 2000.
The FOT generated Power plots for HRDI to aid them in their investigation of
the red limit they see in telemetry. No problems have been identified.
A command sheet for the PEM calibration was implemented on Sunday. The
calibration requires taking them out of PMON control for two orbits and then
returning them to PMON control. All went well.
New Activities Started/ Milestones Approaching
The FOT continued scheduling all available TDRS supports in an effort maximize
real-time data collection. In addition, the FOT continued to use the Pass 2
side of NBTR-B to capture sunrise and sunset data, per the request of HALOE
Operations.
The Excel spreadsheet that generates the statistical data for the power
subsystem has stopped running. The FOT is trying to resolve this problem.
MEETINGS
The FOT did not attend any meetings during this report period.
FUTURE SIGNIFICANT EVENTS
No specific date was agreed upon for UARS Safehold configuration discussions.
Instead, the Supporting Engineers involved will be contacted every two weeks to
determine if either the Flight Operations Team or Supporting Engineers have any
new developments that should be discussed in a group.
The next UARS Yaw slew is scheduled for Febuary 15, 2000 (DOY 006). This
maneuver will be a Forward to Reverse Flight maneuver (Northern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 7 January 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
45413 002/0708 00-001 NBTR NBTR NOOPS
45421 002/1939 00-002 NBTR NBTR NOOPS
45439 003/2317 00-003 HALOE Power OFF
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 7 January 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ------------ ------
HRPTSPBIMON R-LO HRDI PTS Pulse Bus Current Mon. 45517
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
No Alerts to report this report period.
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 7 January 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- ----------------
None this report period
______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 7 January 2000
Not Available
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
------ ---- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 7 January 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-001 FOT See list January 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-089 FOT See list December 14170 Generic Late Acquisition
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiving 512 PB
SEVNTFW00/01